a material having and index of refraction of 1.30 is used as an anti-reflective coating on a piece of glass (n=1.50). what should the minimum thickness of the thin film be to minimize reflection of 500-nm light
Answer #1
Concepts and reason

This question is based upon the concept of the destructive interference.

Firstly, write the equation for destructive interference. Rearrange the equation for thickness and calculate the value of minimum thickness of the thin film. Destructive interference occurs

Fundamentals

The expression for destructive interference is,

2nt=(m+12)λ2nt = \left( {m + \frac{1}{2}} \right)\lambda

Here, λ\lambda is the wavelength, mm is the order number, tt is the thickness of film and nn is the index of the refraction.

The expression for destructive interference is,

2nt=(m+12)λ2nt = \left( {m + \frac{1}{2}} \right)\lambda

Rearrange the equation for thickness of non-reflecting film .

2nt=(m+12)λt=(m+12)λ2n\begin{array}{c}\\2nt = \left( {m + \frac{1}{2}} \right)\lambda \\\\t = \frac{{\left( {m + \frac{1}{2}} \right)\lambda }}{{2n}}\\\end{array}

Minimum value of mm is zero.

Substitute 00 for mm in the above equation.

t=(0+12)λ2nt=λ4n\begin{array}{c}\\t = \frac{{\left( {0 + \frac{1}{2}} \right)\lambda }}{{2n}}\\\\t = \frac{\lambda }{{4n}}\\\end{array}

The expression for thickness of non-reflecting film is,

t=λ4nt = \frac{\lambda }{{4n}}

Substitute 500nm500{\rm{ nm}} for λ\lambda and 1.31.3 for nn in above equation of height.

t=500nm4(1.3)=96nm\begin{array}{c}\\t = \frac{{500{\rm{ nm}}}}{{4\left( {1.3} \right)}}\\\\ = 96{\rm{ nm}}\\\end{array}

Ans:

The value of thickness of non-reflecting film is 96nm96{\rm{ nm}} .